Discussion

authorname, affiliation author2name, affiliation

If you have sections, use this style

Reference style

Here are two examples of the reference style we are using. Experimental evidence supporting the first hypothesis has been provided by several authors (Rudolph et al., 1967; Heyman et al., 1977; von Ritter et al., 1988). For a precise description of these models we refer to Almeida and Gidas (1993).

Figures

Figures can be included as .ps files. Psfig is an efficient way of handling postscript files. There are many sources of info about Psfig on the World Wide Web. If you do not have the psfig.sty file available to you, please let me know and I will email it to you.

Final Steps

When your paper is in final form, please email me the L^{A}TEX file (don’t forget any included .ps files) and send a hard copy. My mailing address is Heidi Sestrich, Dept. of Statistics, Baker Hall 232A, Carnegie Mellon University, Pittsburgh, PA 15213, email is heidi@stat.cmu.edu.

Remember that we are preparing a camera ready copy for Springer-Verlag. The published paper will look *exactly* the same as the paper you send me.

References

Almeida, M. and Gidas, B. (1993). A Variational Method for Estimating the Parameters of MRF from Complete or Incomplete Data, Annals Appl. Prob. 3, 103-136.

Heyman, M.A., Payne B.D., Hoffman, J.I.E. and Rudolph, A.M. (1977). Blood flow measurements with radionuclide-labled particles. Progress in Cardiovascular Diseases 10, 55-79.